### v1

# Normalized solutions to mixed dispersion nonlinear Schrödinger system with coupled nonlinearity

Abstract
In this paper, we consider the existence of normalized solutions for the following biharmonic nonlinear Schrödinger system
$\begin{cases}\Delta^{2}u+\alpha_{1}\Delta u+\lambda u=\beta r_{1}|u|^{r_{1}-2}|v|^{r_{2}}u&\text{ in }\mathbb{R}^{N},\\ \Delta^{2}v+\alpha_{2}\Delta v+\lambda v=\beta r_{2}|u|^{r_{1}}|v|^{r_{2}-2}v&\text{ in }\mathbb{R}^{N},\\ \int_{\mathbb{R}^{N}}(u^{2}+v^{2})\mathrm{d}x=\rho^{2},\end{cases}$
where $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 1$. $\rho^{2}$ stands for the prescribed mass, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier. Such single constraint permits mass transformation in two materials. When $r_{1}+r_{2}\in\left(2,2+\frac{8}{N}\right]$, we obtain a dichotomy result for the existence of nontrivial ground states. Especially when $\alpha_{1}=\alpha_{2}$, the ground state exists for all $\rho>0$ if and only if $r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}$. When $r_{1}+r_{2}\in\left(2+\frac{8}{N},\frac{2N}{(N-4)^{+}}\right)$ and $N\geq 2$, we obtain the existence of radial nontrivial mountain pass solution for small $\rho>0$.
Keywords: Biharmonic system, Normalized solutions, Ground state, Mountain pass solution.
1 Introduction
This paper is concerned with the existence of solutions for the biharmonic nonlinear Schrödinger system
$\begin{cases}\Delta^{2}u+\alpha_{1}\Delta u+\lambda u=\beta r_{1}|u|^{r_{1}-2}|v|^{r_{2}}u&\text{ in }\mathbb{R}^{N},\\ \Delta^{2}v+\alpha_{2}\Delta v+\lambda v=\beta r_{2}|u|^{r_{1}}|v|^{r_{2}-2}v&\text{ in }\mathbb{R}^{N},\end{cases}$ (1.1)
under the mass constraint
$\int_{\mathbb{R}^{N}}(u^{2}+v^{2})\mathrm{d}x=\rho^{2},$ (1.2)
where $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $N\geq 1$, $r_{1}$, $r_{2}>1$, $r:=r_{1}+r_{2}\in(2,2^{**})$, $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier, which is unknown. Here
$2^{**}:=\frac{2N}{(N-4)^{+}},\,\mbox{namely $2^{**}=\frac{2N}{N-4}$ if $N\geq 5$, and $2^{**}=+\infty$ if $N\leq 4$},$
is called Sobolev critical exponent. Any solution $(u,v,\lambda)$ of (1.1) satisfying (1.2) is usually called normalized solution.
In recent years, many researchers have considered the normalized solutions for the following single biharmonic equation
$\begin{cases}\Delta^{2}u+\alpha\Delta u+\lambda u=f(u)&\text{ in }\mathbb{R}^{N},\\ \int_{\mathbb{R}^{N}}u^{2}\mathrm{d}x=\rho^{2},\end{cases}$ (1.3)
where $\alpha\in\mathbb{R}$, $\rho>0$ and $\lambda$ is a Lagrange multiplier. In order to regularize and stabilize the solutions to the Schrödinger equation
$i\partial_{t}\psi+\Delta\psi+f(\psi)=0\quad\mbox{ in }\mathbb{R}^{N}\times(0,\infty),$
Karpman and Shagalov (see [11] and references therein) proposed the fourth order equation
$i\partial_{t}\psi-\gamma\Delta^{2}\psi+\beta\Delta\psi+f(\psi)=0\quad\mbox{ in}\mathbb{R}^{N}\times(0,\infty).$ (1.4)
On the other hand, to prevent blow-up in finite time, Fibich, Ilan and Papanicolaou [9] also added a small fourth-order dispersion term in nonlinear Schrödinger equation as a nonparaxial correction. For more introduction about the background of (1.4), we refer to [1, 2].
Here we are concerned with the standing wave solutions, namely $\psi$ of the form
$\psi(x,t)=e^{i\lambda t}u(x).$ (1.5)
If $f(\psi)=e^{i\lambda t}f(u)$, up to a scaling, (1.4)-(1.5) with prescribed $L^{2}$-norm can be reduced to the equation (1.3). Note that the biharmonic term and Laplacian term are two dispersive terms, and a lack of homogeneity occurs provided $\alpha\neq 0$, which brings main difficulties in deriving the existence of solutions. One of approaches to obtain solutions of (1.3) is variational method. That is, the critical points of the associated energy functional
$E(u)=\frac{1}{2}\|\Delta u\|_{2}^{2}-\frac{\alpha}{2}\|\nabla u\|_{2}^{2}-\int_{\mathbb{R}^{N}}\int_{0}^{u}f(t)\mathrm{d}t\mathrm{d}x,\quad\forall u\in\Lambda_{\rho},$
correspond to the solutions of (1.3), where the constraint
$\Lambda_{\rho}:=\{u\in H^{2}(\mathbb{R}^{N}):\|u\|_{2}^{2}=\rho^{2}\},$
and the Sobolev space $H^{2}(\mathbb{R}^{N})$ is defined as follows
$H^{2}(\mathbb{R}^{N}):=\left\{u\in L^{2}(\mathbb{R}^{N}):\nabla u,\Delta u\in L^{2}(\mathbb{R}^{N})\right\}$
endowed with the equivalent norm $\|u\|:=\left(\|\Delta u\|^{2}_{2}+\|u\|^{2}_{2}\right)^{\frac{1}{2}}$. Observe that the biharmonic term possesses dominating role in energy functional when $\alpha\leq 0$, while the Laplacian term has great effect on the geometry of energy functional when $\alpha>0$, especially which shows a convex-concave shape when we rescale the function $u$. Regardless of $\alpha\leq 0$ or $\alpha>0$, there exist a lot of works concentrated on the nonlinearities $f(u)=\mu|u|^{q-2}u+|u|^{p-2}u$ with $\mu\geq 0$.
The simplest case is the power nonlinearities $f(u)=|u|^{p-2}u$. Here we are convenient to introduce the parameter $\bar{r}:=2+\frac{8}{N}$, which is usually called $L^{2}$-critical exponent or mass critical exponent for fourth order equation. By Gagliardo-Nirenberg inequality (see Lemma 2.1 below) or dilations, the sign of $p-\bar{r}$ decides the geometry of the energy functional $E$ on $\Lambda_{\rho}$. Many researchers treated this type of problem according to the value of $p$ and the sign of $\alpha$.
  • When $\alpha\leq 0$, Bonheure et al. [1] considered the coercive case, namely $p\in(2,\bar{r})$. Observe that the minimization level
    $\overline{m}(\rho):=\underset{u\in\Lambda_{\rho}}{\inf}E(u)$
    is obviously sub-additive since the energy is invariant under translation. Whenever $\overline{m}(\rho)$ is negative, the vanishing of the minimizing sequence will not occur, hence $\overline{m}(\rho)$ can be attained. They proved that $\overline{m}(\rho)<0$ if $2<p<2+\frac{4}{N}$, so $\overline{m}(\rho)$ is attainable for all $\rho>0$. For $2+\frac{4}{N}\leq p<\bar{r}$, $\overline{m}(\rho)$ could be zero if the mass is large. As a result, the existence is a dichotomy result with respect to mass.
  • Bonheure et al. [2] investigated the case $\alpha<0$ and $p\in[\bar{r},2^{**})$. They proved the existence of the ground states when $\rho\in(c_{1},c_{2})$ for two numbers $c_{1}\geq 0,\,c_{2}\in(0,\infty]$ depending only on $N,\,p$. In particular, there holds that $c_{1}=0$ if $p>\bar{r}$. Moreover, they constructed minimax levels by $\mathbb{Z}_{2}$-genus, obtaining the multiplicity of radial normalized solutions.
  • When $\alpha>0$, the problem is more involved since the Laplacian has much effect on the shape of energy functional. Fernández et al. [8] established some non-homogeneous Gagliardo-Nirenberg inequalities, and showed some overall results about the existence and non-existence of global minimizers for $p\in(2,\bar{r}]$ and local minimizers for $p\in(\bar{r},2^{**})$. Furthermore, when $p\in(\bar{r},2^{**})$, Luo and Yang [14] proved that (1.3) has another mountain pass type solution as $\rho$ is small.
Next we introduce some recent results about $f(u)=\mu|u|^{q-2}u+|u|^{p-2}u$ with $\mu>0$.
  • When $\alpha=0$, $N\geq 5$ and $p=2^{**}$, Ma and Chang [16] showed that for $q\in(2,\bar{r})$, (1.3) has a ground state solution provided $\rho$ is small; Liu and Zhang [13] proved that for $q\in(2,\bar{r})$, there exists a mountain pass type solution of (1.3) for small $\rho$.
  • When $\alpha=0$, $N\geq 5$, $\bar{r}\leq q<p\leq 2^{**}$, Chang et al. [5] discussed the existence, non-existence of normalized solutions for (1.3), and they also proved the strong instability of standing waves.
Finally, for more general $L^{2}$-subcritical nonlinearities and $\alpha\in\mathbb{R}$, Luo and Zhang [15] proved a dichotomy result with respect to the mass for the existence of ground states to (1.3). Chen and Chen [6] considered nonlinearities $f$ involving Hardy-Littlewood-Sobolev upper critical and combined nonlinearities.
As far as we are aware, it seems that there is no paper considering normalized solutions involving mixed dispersion nonlinear Schrödinger system. Motivated by above works, we are interested in the existence of normalized solutions for problem (1.1)-(1.2). That is, we search for $(u,v,\lambda)$ satisfying (1.1)-(1.2). A solution $(u,v,\lambda)$ is said nontrivial, which means that $u\neq 0$ and $v\neq 0$. The functional
$I(u,v)=\frac{1}{2}\int_{\mathbb{R}^{N}}(|\Delta u|^{2}+|\Delta v|^{2})\mathrm{d}x-\frac{\alpha_{1}}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\mathrm{d}x-\frac{\alpha_{2}}{2}\int_{\mathbb{R}^{N}}|\nabla v|^{2}\mathrm{d}x-\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x$
is the corresponding variational functional of problem (1.1)-(1.2) defined on the constraint
$S_{\rho}:=\left\{(u,v)\in H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}):\|u\|^{2}_{2}+\|v\|^{2}_{2}=\rho^{2}\right\}.$ (1.6)
To the best of our knowledge, very few works concerned the single constraint for $u$ and $v$ in nonlinear Schrödinger system. However the constraint (1.6) could be often encountered in physical world. In fact, $(u,v)\in S_{\rho}$ represents that the total mass of system is conserved, but the mass of $u$ and $v$ may transform mutually, where the transformation may be aroused due to chemical reactions or the movement in physical space.
Note that the energy functional $I$ under the different cases $r\in(2,\bar{r})$, $r=\bar{r}$ and $r\in(\bar{r},2^{**})$ has different geometry on $S_{\rho}$. When either $r\in(2,\bar{r})$ or $r=\bar{r}$ and $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ ($\mathcal{D}_{1}$ is given in Theorem 1.1 below), we know that $I$ is coercive on $S_{\rho}$, see Lemma 3.1. Hence we concern whether the minimization problem
$m(\rho):=\inf_{(u,v)\in S_{\rho}}I(u,v)$ (1.7)
is attainable, whose minimizers are usually called ground states. As the method of [8, 15], the key ingredient is to establish the compactness of minimizing sequence. We emphasize here that $\alpha_{1}$, $\alpha_{2}>0$ and the fact that we can not obtain the strong convergence $\nabla u_{n}\rightarrow\nabla u$ in $L^{2}(\mathbb{R}^{N})$ only from $u_{n}\rightharpoonup u$ in $H^{2}(\mathbb{R}^{N})$ bring the main difficulties for problem (1.1)-(1.2). For that, we first establish a strict sub-additivity for $m(\rho)$ in Lemma 3.3. With this result and Lions’ concentration compactness principle, we provide in Theorem 3.5 an alternative result for minimizing sequence, that is, the minimizing sequence either vanishes or has a convergent subsequence.
On the other hand, we consider another auxiliary minimization problem
$m^{J}(\rho):=\underset{(u,v)\in S_{\rho}}{\inf}J(u,v),$ (1.8)
where
$J(u,v)=\frac{1}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{1}{2}\alpha_{1}\|\nabla u\|^{2}_{2}-\frac{1}{2}\alpha_{2}\|\nabla v\|^{2}_{2},\quad\forall(u,v)\in S_{\rho}.$
It can be seen that $J$ is coercive on $S_{\rho}$, so $m^{J}(\rho)$ is always well-defined. We will show in Lemma 3.2 that $m^{J}(\rho)$ is never achieved. If the minimizing sequence for $m(\rho)$ vanishes, there must hold $m(\rho)=m^{J}(\rho)$. Therefore we can rule out the vanishing of the minimizing sequence if $m(\rho)<m^{J}(\rho)$. So the key step is to establish the comparison between $m(\rho)$ and $m^{J}(\rho)$. As the ideas in [8], this comparison can be reduced into whether the supremum of
$Q(u,v):=\frac{\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x}{\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\|v\|_{2}^{2}\right)\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{r}{2}-1}}$ (1.9)
in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ is finite, see subsection 3.1. After careful analysis, we obtain the following dichotomy result.
Theorem 1.1.
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$. Let $\bar{r}=2+\frac{8}{N}$, $r=r_{1}+r_{2}$ and $\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r}$, where ${C}_{N,r}$ is the optimal constant for Gagliardo-Nirenberg inequality (2.1). Then there exists some
$\rho^{*}\in\begin{cases}[0,\infty)\quad&\mbox{if}\;\;r<\bar{r},\\ \left[0,\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r},\end{cases}$
such that $m(\rho)$ can be attained if
$\rho\in\begin{cases}(\rho^{*},\infty)\quad&\mbox{if}\;\;r<\bar{r},\\ \left(\rho^{*},\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r},\end{cases}$
where the minimizers are nontrivial. Moreover, (1.7) can never be attained if $\rho<\rho*$.
Remark 1.2.
From the proof of Theorem 1.1, the strict subadditivity is important for obtaining the existence of solutions. While considering
$\begin{cases}\Delta^{2}u+\alpha_{1}\Delta u+\lambda_{1}u=\beta r_{1}|u|^{r_{1}-2}|v|^{r_{2}}u&\text{ in }\mathbb{R}^{N},\\ \Delta^{2}v+\alpha_{2}\Delta v+\lambda_{2}v=\beta r_{2}|u|^{r_{1}}|v|^{r_{2}-2}v&\text{ in }\mathbb{R}^{N},\end{cases}$
with two constraints
$\int_{\mathbb{R}^{N}}u^{2}\mathrm{d}x=\rho^{2}_{1},\leavevmode\nobreak\ \int_{\mathbb{R}^{N}}v^{2}\mathrm{d}x=\rho^{2}_{2}\quad\text{with}\leavevmode\nobreak\ \rho_{1},\leavevmode\nobreak\ \rho_{2}>0,$
the existence issue is very different from single constraint case, since it could not establish the strict subadditivity if we proceed as the proof of Lemma 3.3.
Notice that $\rho^{*}$ in Theorem 1.1 has chance to be 0, this means that $m(\rho)$ can be achieved for all $\rho>0$. Next, we aim to find the borderline to guarantee $\rho^{*}=0$. To this end, the cases that $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$ have great differences. Actually the problem in the case $\alpha_{1}=\alpha_{2}$ is very similar to the single equation
$\begin{cases}\Delta^{2}u+\alpha_{1}\Delta u+\lambda u=\beta|u|^{r-2}u&\text{ in }\mathbb{R}^{N},\\ \int_{\mathbb{R}^{N}}u^{2}\mathrm{d}x=\rho^{2}.\end{cases}$
We will in subsection 3.2 make use of the estimates of [8] to give a borderline for $\rho^{*}=0$. However, when $\alpha_{1}\neq\alpha_{2}$, the estimates are challenging and elusive. We only know that the same condition as the case $\alpha_{1}=\alpha_{2}$ can guarantee $\rho^{*}>0$, but it is difficult for us to conclude how $\rho^{*}$ equals 0. Our result is the following.
Theorem 1.3.
Assume $\alpha_{1},\,\alpha_{2},\,\beta>0$ and $r\leq\bar{r}$. Let $\rho^{*}$ be given in Theorem 1.1. Then $\rho^{*}>0$ if
$\max\left\{4,2+\frac{8}{N+1}\right\}\leq r\leq 2+\frac{8}{N}.$ (1.10)
In particular, when $\alpha_{1}=\alpha_{2}$, (1.10) is also a necessary condition for $\rho^{*}>0$.
Remark 1.4.
Theorems 1.1 and 1.3 can tell us that when $\alpha_{1}=\alpha_{2}$, the nontrivial ground state exists for all $\rho>0$ if and only if
$2<r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}.$
Now, we consider the case $r\in(\bar{r},2^{**})$. Observe that when the mass is small, the energy functional $I$ possesses a mountain pass geometry. More precisely, when $\beta\rho^{r-2}<c^{*}(N,r)$, there exist $R_{1}>R_{0}>0$ such that for any $c\in(R_{0},R_{1})$, $I$ has a positive lower bound on $S_{\rho}$ with the restrict $\|\Delta u\|_{2}^{2}+\|\Delta v\|_{2}^{2}=c^{2}$, where
$c^{*}(N,r):=\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r\gamma_{r}-2}{(r\gamma_{r}-1)\max\{\alpha_{1},\alpha_{2}\}}\right)^{r\gamma_{r}-2},$
$\gamma_{r}:=\frac{N(r-2)}{4r}$ and $\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r}$. On the other hand, we can find two points $(u_{0},v_{0})$, $(u_{1},v_{1})\in S_{\rho}$ such that $I(u_{0},v_{0}),\,I(u_{1},v_{1})<0$ with $\|\Delta u_{0}\|_{2}^{2}+\|\Delta v_{0}\|_{2}^{2}<R_{0}^{2}$ and $\|\Delta u_{1}\|_{2}^{2}+\|\Delta v_{1}\|_{2}^{2}>R_{1}^{2}$. Thus, we can establish a mountain pass structure, precisely see Section 4. Using Jeanjean’s method in [10], we can obtain a Palais-Smale sequence $\{(u_{n},v_{n})\}$ approaching Pohožaev manifold, but whose compactness is very involved. By virtue of Lagrange multiplier rule, we derive a sequence of Lagrange multipliers $\{\lambda_{n}\}$ corresponding to the sequence $\{(u_{n},v_{n})\}$. For the compactness, a key step is to deduce $\underset{n\to\infty}{\liminf}\,\lambda_{n}>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. Following the method in [14], if in addition $\beta\rho^{r-2}<c_{*}(N,r)$, we can reach this aim, where
$c_{*}(N,r)$ :=$&\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{1-\gamma_{r}}{\gamma_{r}}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\quad\quad\quad\text{ if }\frac{1}{2}<\gamma_{r}<1,\\ &\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r-2}{2(r\gamma_{r}-1)}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\leavevmode\nobreak\ \quad\text{ if }0<\gamma_{r}\leq\frac{1}{2}.\end{aligned}\right.$
Our result for $r\in(\bar{r},2^{**})$ can be stated as follows.
Theorem 1.5.
Let $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 2$ and $r:=r_{1}+r_{2}\in(\bar{r},2^{**})$. Then for any $\rho>0$ satisfying
$\beta\rho^{r-2}<\min\left\{c^{*}(N,r),c_{*}(N,r)\right\},$
problem (1.1)-(1.2) has a mountain pass type nontrivial radial solution for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
Remark 1.6.
Here we assume $N\geq 2$, because in Section 4, we will replace $H^{2}(\mathbb{R}^{N})$ by the radial subspace
$H_{r}^{2}(\mathbb{R}^{N}):=\{u\in H^{2}(\mathbb{R}^{N}):u\mbox{ is radially symmetric}\}.$
The compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ for $2<p<2^{**}$ only hold when $N\geq 2$.
This paper is organized as follows. In Section 2, we present some useful lemmas. In Section 3, we give the proof of Theorems 1.1 and 1.3. In Section 4, we give the proof of Theorem 1.5. In the rest of this paper, unless otherwise specified, we always assume that $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $r_{1}$, $r_{2}>1$.
2 Preliminaries
In this section, we are devoted to some notations and preliminary results.
Lemma 2.1 (Gagliardo-Nirenberg inequality [17]).
For $N\geq 1$ and $2<p<2^{**}$, there exists an optimal constant $C_{N,p}>0$ depending on $N$, $p$ such that
$\|u\|_{p}\leq C_{N,p}\|u\|_{2}^{1-\gamma_{p}}\|\Delta u\|_{2}^{\gamma_{p}},\quad\forall u\in H^{2}(\mathbb{R}^{N}).$ (2.1)
Applying classical Fourier transform and Hölder inequality, we can easily get the interpolation inequality
$\|\nabla u\|^{2}_{2}\leq\|u\|_{2}\cdot\|\Delta u\|_{2},\quad\forall u\in H^{2}(\mathbb{R}^{N}).$ (2.2)
Using Hölder’s inequality, Lemma 2.1 and Young’s inequality, we have
$\int|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x$ $\leq\left(\int\left|u\right|^{r}\right)^{\frac{r_{1}}{r}}\left(\int\left|v\right|^{r}\right)^{\frac{r_{2}}{r}}$
$\leq C_{N,r}^{r}\left(\|u\|_{2}^{r_{1}}\|v\|_{2}^{r_{2}}\right)^{1-\gamma_{r}}\left(\|\Delta u\|_{2}^{r_{1}}\|\Delta v\|_{2}^{r_{2}}\right)^{\gamma_{r}}$ (2.3)
$\leq\mathcal{D}_{1}\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{r\left(1-\gamma_{r}\right)}{2}}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)^{\frac{r\gamma_{r}}{2}},$
where $\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r}$. In the last inequality above, we used the basic inequality
$a^{r_{1}}b^{r_{2}}\leq\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}\left(a^{2}+b^{2}\right)^{\frac{r}{2}},\quad\forall a,b>0,$
where the equality holds if and only if $\frac{a^{2}}{b^{2}}=\frac{r_{1}}{r_{2}}$. As a result, we obtain the following lemma.
Lemma 2.2.
The equalities in (2) hold if and only if $v=\sqrt{\frac{r_{2}}{r_{1}}}u$ and $u$ is an extremal for Gagliardo-Nirenberg inequality given in (2.1).
By (2.2) and Cauchy inequality, we deduce that
$\alpha_{1}\|\nabla u\|^{2}_{2}+\alpha_{2}\|\nabla v\|^{2}_{2}\leq$ $\alpha_{1}\|u\|_{2}\|\Delta u\|_{2}+\alpha_{2}\|v\|_{2}\|\Delta v\|_{2}$ (2.4)
$\leq$ $\mathcal{D}_{2}\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{1}{2}}(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2})^{\frac{1}{2}},$
where $\mathcal{D}_{2}:=\max\{\alpha_{1},\alpha_{2}\}$.
As the proof of [3, Lemma 2.1] and [8, Remark 3.10], we give the Pohožaev identity for the problem (1.1) and omit it’s proof.
Lemma 2.3.
Assume that $(u,v)\in H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ solves (1.1), then the Pohožaev identity holds:
$\frac{N-4}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{N-2}{2}\left(\alpha_{1}\|\nabla u\|^{2}_{2}+\alpha_{2}\|\nabla v\|^{2}_{2}\right)+\frac{N}{2}\lambda\left(\|u\|^{2}_{2}+\|v\|^{2}_{2}\right)$ (2.5)
$=$ $N\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
Lemma 2.4.
Let $\alpha_{1},\,\alpha_{2},\,\beta>0$ and $r\in(2,2^{**}]$. When $\lambda\leq 0$, then problem (1.1)-(1.2) admits no solutions.
Proof.
Assume that $(u,v)$ is a solution of (1.1)-(1.2). Multiplying the equation (1.1) by $u$ and $v$ respectively, we get that
$\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}-\alpha_{1}\|\nabla u\|^{2}_{2}-\alpha_{2}\|\nabla v\|^{2}_{2}+\lambda\|u\|^{2}_{2}+\lambda\|v\|^{2}_{2}=\beta r\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$ (2.6)
By (2.5) and (2.6), we have
$-\alpha_{1}\|\nabla u\|^{2}_{2}-\alpha_{2}\|\nabla v\|^{2}_{2}+2\lambda(\|u\|^{2}_{2}+\|v\|^{2}_{2})=\beta\left(N-r\frac{N-4}{2}\right)\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$ (2.7)
Since $\alpha_{1},\,\alpha_{2},\,\beta>0$ and $\lambda\leq 0$, when $r\in(2,2^{**}]$ and by (2.7), we can deduce that $\nabla u$, $\nabla v=0$ a.e. in $\mathbb{R}^{N}$, which contradicts $(u,v)\in S_{\rho}$. ∎
3 Mass-subcritical and mass-critical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r=r_{1}+r_{2}\in(2,\bar{r}]$. Initially, we obtain some important properties about $m(\rho)$ given in (1.7).
Lemma 3.1.
Let $\rho>0$, $r\in(2,\bar{r}]$ and $\mathcal{D}_{1}$ be given in (2).
$\mathrm{(i)}$ $m(\rho)$ is finite if and only if either $r\in(2,\bar{r})$ or $r=\bar{r}$ and $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. Moreover, in case $r=\bar{r}$ and $\rho\geq\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$, we have $m(\rho)=-\infty$.
$\mathrm{(ii)}$ If either $r\in(2,\bar{r})$ or $r=\bar{r}$ and $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$, then the map $\rho\mapsto m(\rho)$ is continuous.
Proof.
(i) By (2) and (2.4), we get that for $(u,v)\in S_{\rho}$,
$I(u,v)=$ $\frac{1}{2}\|\Delta u\|^{2}_{2}+\frac{1}{2}\|\Delta v\|^{2}_{2}-\frac{\alpha_{1}}{2}\|\nabla u\|^{2}_{2}-\frac{\alpha_{2}}{2}\|\nabla v\|^{2}_{2}-\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x$ (3.1)
$\geq$ $\frac{1}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{1}{2}\mathcal{D}_{2}\rho\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)^{\frac{1}{2}}$
$-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)^{\frac{r\gamma_{r}}{2}}.$
When $r\in(2,\bar{r})$, we get that $r\gamma_{r}<2$ and $I$ is coercive on $S_{\rho}$, which implies that $m(\rho)>-\infty$.
When $r=\bar{r}$ and $\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}<\frac{1}{2}$, that is $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$, we have $r\gamma_{r}=2$ and $I$ is also coercive on $S_{\rho}$. Still $m(\rho)>-\infty$. If $\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}\geq\frac{1}{2}$, we claim $I$ is unbounded from below on $S_{\rho}$. Let $U$ be an extremal of Gagliardo-Nirenberg inequality given in (2.1) with $p=\bar{r}$ and $\|U\|_{2}^{2}=\frac{\rho^{2}r_{1}}{r}$. As we remarked in Lemma 2.2, the equalities in (2) can hold if $u=U$ and $v=\sqrt{\frac{r_{2}}{r_{1}}}U$. Setting $U_{t}:=t^{\frac{N}{2}}U(tx)$, $u_{t}:=t^{\frac{N}{2}}u(tx)$ and $v_{t}:=t^{\frac{N}{2}}v(tx)$, then $\|U_{t}\|_{2}^{2}=\|U\|_{2}^{2}$ and
$I(u_{t},v_{t})=$ $\left(\frac{1}{2}-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}\right)\frac{r}{r_{1}}\|\Delta U_{t}\|^{2}_{2}-\frac{\alpha_{1}}{2}\|\nabla U_{t}\|^{2}_{2}-\frac{\alpha_{2}}{2}\frac{r_{2}}{r_{1}}\|\nabla U_{t}\|^{2}_{2}$
$\leq$ $-\frac{\alpha_{1}t^{2}}{2}\|\nabla U\|^{2}_{2}-\frac{\alpha_{2}t^{2}}{2}\frac{r_{2}}{r_{1}}\|\nabla U\|^{2}_{2}.$
Thus, letting $t\to\infty$ in the above, we get the desired conclusion.
(ii) Let $\rho_{n}\rightarrow\rho$ as $n\rightarrow\infty$, and in addition $\rho,\rho_{n}<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ if $r=\bar{r}$. There exists $(u_{n},v_{n})\in S_{\rho_{n}}$ such that
$m(\rho_{n})\leq I(u_{n},v_{n})<m\left(\rho_{n}\right)+\frac{1}{n}.$ (3.2)
By the coerciveness of $I$ on $S_{\rho}$, $\rho_{n}\to\rho$ and (3.2), we can easily deduce that $\left\{(u_{n},v_{n})\right\}$ is bounded in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$. Let $\widetilde{u}_{n}:=\frac{\rho}{\rho_{n}}u_{n}$ and $\widetilde{v}_{n}:=\frac{\rho}{\rho_{n}}v_{n}$. Then $\{(\widetilde{u}_{n},\widetilde{v}_{n})\}\subset S_{\rho}$ and
$m(\rho)\leq$ $I(\widetilde{u}_{n},\widetilde{v}_{n})$
$=$ $I(u_{n},v_{n})+\frac{1}{2}\left(\frac{\rho^{2}}{\rho_{n}^{2}}-1\right)\int_{\mathbb{R}^{N}}(|\Delta u_{n}|^{2}+|\Delta v_{n}|^{2})\mathrm{d}x-\frac{\alpha_{1}}{2}\left(\frac{\rho^{2}}{\rho_{n}^{2}}-1\right)\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}\mathrm{d}x$
$-\frac{\alpha_{2}}{2}\left(\frac{\rho^{2}}{\rho_{n}^{2}}-1\right)\int_{\mathbb{R}^{N}}|\nabla v_{n}|^{2}\mathrm{d}x-\beta\left(\frac{\rho^{r}}{\rho_{n}^{r}}-1\right)\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x$
$=$ $I(u_{n},v_{n})+o_{n}(1)$
due to $\rho_{n}\rightarrow\rho$. Thus
$m(\rho)\leq\liminf_{n\rightarrow\infty}m\left(\rho_{n}\right).$ (3.3)
On the other hand, define $\widetilde{w}^{1}_{n}:=\frac{\rho_{n}}{\rho}w^{1}_{n}$ and $\widetilde{w}^{2}_{n}:=\frac{\rho_{n}}{\rho}w^{2}_{n}$, where $\left\{(w^{1}_{n},w^{2}_{n})\right\}\subset S_{\rho}$ is a minimizing sequence for $m(\rho)$. By a similar discussion as above, $\left\{(w^{1}_{n},w^{2}_{n})\right\}$ is bounded in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$, and
$m\left(\rho_{n}\right)\leq I\left(\widetilde{w}^{1}_{n},\widetilde{w}^{2}_{n}\right)=I\left(w^{1}_{n},w^{2}_{n}\right)+o_{n}(1)=m(\rho)+o_{n}(1).$
Thus,
$\limsup_{n\rightarrow\infty}m\left(\rho_{n}\right)\leq m(\rho),$ (3.4)
Combining (3.3) with (3.4), we obtain that $m(\rho)=\lim\limits_{n\rightarrow\infty}m(\rho_{n})$. Thus (ii) holds. ∎
Let $m^{J}(\rho)$ be given in (1.8), which is well-defined for all $\rho>0$. The following result can give the accurate value of $m^{J}(\rho)$, and it is not attainable.
Lemma 3.2.
For any $\rho>0$, we have $m^{J}(\rho)=-\frac{\max\{\alpha_{1}^{2},\alpha_{2}^{2}\}}{8}\rho^{2}$, which is never achieved.
Proof.
By (1.8), we can rewrite
$m^{J}(\rho)=\underset{\rho_{1}^{2}+\rho_{2}^{2}=\rho^{2}}{\inf}M(\rho_{1},\rho_{2})$ (3.5)
where
$M(\rho_{1},\rho_{2}):=\underset{u,v\in H^{2}(\mathbb{R}^{N}),\,\|u\|_{2}^{2}=\rho_{1}^{2},\,\|v\|_{2}^{2}=\rho_{2}^{2}}{\inf}J(u,v).$
It can easily see that
$M(\rho_{1},\rho_{2})=$ $\underset{u\in H^{2}(\mathbb{R}^{N}),\,\|u\|_{2}^{2}=\rho_{1}^{2}}{\inf}\left(\frac{1}{2}\|\Delta u\|^{2}_{2}-\frac{\alpha_{1}}{2}\|\nabla u\|^{2}_{2}\right)$
$+\underset{v\in H^{2}(\mathbb{R}^{N}),\,\|v\|_{2}^{2}=\rho_{2}^{2}}{\inf}\left(\frac{1}{2}\|\Delta v\|^{2}_{2}-\frac{\alpha_{2}}{2}\|\nabla v\|^{2}_{2}\right).$
As a result, it follows from [4, Lemma 3.1] that
$M(\rho_{1},\rho_{2})=-\frac{\alpha_{1}^{2}}{8}\rho_{1}^{2}-\frac{\alpha_{2}^{2}}{8}\rho_{2}^{2}.$ (3.6)
By (3.5) and (3.6), we have
$m^{J}(\rho)=-\frac{\max\{\alpha_{1}^{2},\alpha_{2}^{2}\}}{8}\rho^{2}.$
Now, we shall prove that $m^{J}(\rho)$ is never achieved. When $\alpha_{1}\neq\alpha_{2}$, we can simply suppose $\alpha_{1}>\alpha_{2}$. At this time, we have $m^{J}(\rho)=-\frac{\alpha_{1}^{2}}{8}\rho^{2}$. Suppose that there exists some $(u,v)\in S_{\rho}$ such that $J(u,v)=-\frac{\alpha_{1}^{2}}{8}\rho^{2}$. We denote $\rho_{1}=\|u\|_{2}$ and $\rho_{2}=\|v\|_{2}$, then
$-\frac{\alpha_{1}^{2}}{8}\rho^{2}=J(u,v)\geq M(\rho_{1},\rho_{2})=-\frac{\alpha_{1}^{2}}{8}\rho_{1}^{2}-\frac{\alpha_{2}^{2}}{8}\rho_{2}^{2}\geq-\frac{\alpha_{1}^{2}}{8}\rho^{2},$
hence $\rho_{1}=\rho$ and $\rho_{2}=0$. So
$\frac{1}{2}\|\Delta u\|^{2}_{2}-\frac{\alpha_{1}}{2}\|\nabla u\|^{2}_{2}=-\frac{\alpha_{1}^{2}}{8}\rho^{2},$
and $u$ is a minimizer for
$M(\rho,0)=\underset{u\in H^{2}(\mathbb{R}^{N}),\,\|u\|_{2}^{2}=\rho^{2}}{\inf}\left(\frac{1}{2}\|\Delta u\|^{2}_{2}-\frac{\alpha_{1}}{2}\|\nabla u\|^{2}_{2}\right).$
However we have already know that $M(\rho,0)$ is never achieved, seeing [4, Lemma 3.1] again. This contradiction tells us that $m^{J}(\rho)$ is also never achieved.
Suppose $\alpha_{1}=\alpha_{2}$ and $(u,v)\in S_{\rho}$ is a minimizer of $m^{J}(\rho)$. With loss of generality, we assume $\rho_{1}=\|u\|_{2}\neq 0$. Thus, $u$ is a minimizer of $M(\rho_{1},0)$, which contradicts [4, Lemma 3.1]. ∎
With the above results, we can show the following sub-additive argument.
Lemma 3.3.
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. If $m(\rho)$ can be attained, then
$\mathrm{(i)}$ for any $\theta>1$, we have $m(\theta\rho)<\theta^{2}m(\rho)$;
$\mathrm{(ii)}$ for any $\rho_{1}>0$, we have $m\left(\left(\rho^{2}+\rho^{2}_{1}\right)^{\frac{1}{2}}\right)<m(\rho)+m(\rho_{1})$, here $\rho_{1}<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ if $r=\bar{r}$.
Proof.
Let $(u,v)\in S_{\rho}$ be a global minimizer of $m(\rho)$. We first claim that $u,v\neq 0$. If otherwise, we can simply assume $u\neq 0$ and $v=0$. Hence
$m^{J}(\rho)\leq J(u,v)=I(u,v)=m(\rho).$
On the other hand, in view of the definition of $m^{J}(\rho)$ and $m(\rho)$, it is obvious that $m^{J}(\rho)\geq m(\rho)$. So $(u,v)$ is a minimizer of $m^{J}(\rho)$, which is a contradiction due to Lemma 3.2. Therefore the claim holds.
(i) For $\theta>1$, by $r\in(2,\bar{r}]$, we obtain
$I\left(\theta u,\theta v\right)=\theta^{2}I\left(u,v\right)+\beta\theta^{2}\left(1-\theta^{r-2}\right)\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x<\theta^{2}I\left(u,v\right)=\theta^{2}m(\rho).$ (3.7)
Thus $m(\theta\rho)\leq I\left(\theta u,\theta v\right)<\theta^{2}I\left(u,v\right)=\theta^{2}m(\rho)$. Hence (i) holds.
(ii) Without loss of generality, we assume that $\rho\geq\rho_{1}>0$. Then by (i),
$m\left(\left(\rho^{2}+\rho^{2}_{1}\right)^{1/2}\right)=$ $m\left(\frac{\left(\rho^{2}+\rho^{2}_{1}\right)^{1/2}}{\rho}\rho\right)<\frac{\rho^{2}+\rho^{2}_{1}}{\rho^{2}}m(\rho)$ (3.8)
$=$ $m(\rho)+\frac{\rho^{2}_{1}}{\rho^{2}}m(\rho)\leq m(\rho)+m(\rho_{1}),$
where we used $\frac{\rho^{2}_{1}}{\rho^{2}}m(\rho)\leq m(\rho_{1})$ in last inequality, whose proof is similar to (i), hence we omit it. Thus (ii) holds. ∎
Lemma 3.4.
Assume that $\left\{(u_{n},v_{n})\right\}$ is a bounded sequence in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$, and it satisfies $\lim\limits_{n\rightarrow\infty}\left(\|u_{n}\|^{2}_{2}+\|v_{n}\|^{2}_{2}\right)=\rho^{2}>0$. Let $\rho_{n}=\frac{\rho}{\left(\|u_{n}\|^{2}_{2}+\|v_{n}\|^{2}_{2}\right)^{1/2}}$, and $\widetilde{u}_{n}=\rho_{n}u_{n}$, $\widetilde{v}_{n}=\rho_{n}v_{n}$. Then the following holds:
$(\widetilde{u}_{n},\widetilde{v}_{n})\in S_{\rho},\quad\lim_{n\rightarrow\infty}\rho_{n}=1,\quad\lim_{n\rightarrow\infty}\left|I\left(\widetilde{u}_{n},\widetilde{v}_{n}\right)-I\left(u_{n},v_{n}\right)\right|=0.$
Proof.
Clearly, $(\widetilde{u}_{n},\widetilde{v}_{n})\in S_{\rho}$ and $\lim\limits_{n\rightarrow\infty}\rho_{n}=1$. So, using the boundedness of $\left\{(u_{n},v_{n})\right\}$, we get that as $n\rightarrow\infty$,
$I\left(\widetilde{u}_{n},\widetilde{v}_{n}\right)-I\left(u_{n},v_{n}\right)=$ $\frac{\rho^{2}_{n}-1}{2}\int_{\mathbb{R}^{N}}(|\Delta u_{n}|^{2}+|\Delta v_{n}|^{2})\mathrm{d}x-\frac{\alpha_{1}(\rho^{2}_{n}-1)}{2}\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}\mathrm{d}x$
$-\frac{\alpha_{2}(\rho^{2}_{n}-1)}{2}\int_{\mathbb{R}^{N}}|\nabla v_{n}|^{2}\mathrm{d}x-\beta(\rho^{r}_{n}-1)\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x$
[…]
The proof is done. ∎
Next, we provide an alternative result for minimizing sequence. This will be used to conclude the existence of minimizers.
Theorem 3.5.
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. Let $\{(u_{n},v_{n})\}\subset S_{\rho}$ be a minimizing sequence of $m(\rho)$ with $\rho>0$. Then one of the following holds:
$\mathrm{(i)}$ (Vanishing)
$\limsup_{n\rightarrow\infty}\sup_{y\in\mathbb{R}^{N}}\int_{B_{1}(y)}\left(|u_{n}|^{2}+|v_{n}|^{2}\right)\mathrm{d}x=0.$ (3.9)
$\mathrm{(ii)}$ (Compactness) Up to a subsequence, there exist $(u,v)\in S_{\rho}$ and a sequence $\left\{y_{n}\right\}\subset\mathbb{R}^{N}$ such that
$\left(u_{n}\left(\cdot-y_{n}\right),v_{n}\left(\cdot-y_{n}\right)\right)\rightarrow(u,v)\quad\mbox{in}\;\,H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$
as $n\rightarrow\infty$, and $(u,v)$ is a global minimizer.
Proof.
Let $\{(u_{n},v_{n})\}\subset S_{\rho}$ be a minimizing sequence for $m(\rho)$, while it does not satisfy (i). Thus,
$0<L:=\limsup_{n\rightarrow\infty}\sup_{y\in\mathbb{R}^{N}}\int_{B_{1}(y)}\left(|u_{n}|^{2}+|v_{n}|^{2}\right)\mathrm{d}x\leq\rho^{2},$ (3.10)
and there exists a sequence $\left\{y_{n}\right\}\subset\mathbb{R}^{N}$ such that up to a subsequence,
$L=\lim_{n\rightarrow\infty}\int_{B_{1}(0)}\left(|u_{n}(x-y_{n})|^{2}+|v_{n}(x-y_{n})|^{2}\right)\mathrm{d}x.$ (3.11)
By the proof of Lemma 3.1, we deduce that the sequence $\left\{(u_{n},v_{n})\right\}$ is bounded in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$. So there exist $(u,v)\in H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ and a renamed subsequence of $\{(u_{n},v_{n})\}$ such that
$\left(u_{n}(\cdot-y_{n}),v_{n}(\cdot-y_{n})\right)\rightharpoonup$ $(u,v)\quad\text{ in }H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}),$
$\left(u_{n}(\cdot-y_{n}),v_{n}(\cdot-y_{n})\right)\rightarrow$ $(u,v)\quad\text{ in }L^{p}_{loc}\left(\mathbb{R}^{N}\right)\times L^{p}_{loc}\left(\mathbb{R}^{N}\right)\text{ for }1\leq p<2^{**},$
$\left(u_{n}(\cdot-y_{n}),v_{n}(\cdot-y_{n})\right)\rightarrow$ $(u,v)\quad\text{ a.e.\leavevmode\nobreak\ in }\mathbb{R}^{N}\times\mathbb{R}^{N},$
and (3.10)-(3.11) imply $(u,v)\neq(0,0)$. Now, let
$\widetilde{u}_{n}:=u_{n}\left(\cdot-y_{n}\right)-u,\quad\widetilde{v}_{n}:=v_{n}\left(\cdot-y_{n}\right)-v.$
By weak convergence and Brezis-Lieb type lemma [7, Lemma 2.3], we can obtain that
$\left\|\Delta u_{n}\right\|_{2}^{2}=$ $\left\|\Delta\left(u+\widetilde{u}_{n}\right)\right\|_{2}^{2}=\|\Delta u\|_{2}^{2}+\left\|\Delta\widetilde{u}_{n}\right\|_{2}^{2}+o_{n}(1),$ (3.12)
$\left\|\Delta v_{n}\right\|_{2}^{2}=$ $\left\|\Delta\left(v+\widetilde{v}_{n}\right)\right\|_{2}^{2}=\|\Delta v\|_{2}^{2}+\left\|\Delta\widetilde{v}_{n}\right\|_{2}^{2}+o_{n}(1),$
$\left\|\nabla u_{n}\right\|_{2}^{2}=$ $\left\|\nabla\left(u+\widetilde{u}_{n}\right)\right\|_{2}^{2}=\|\nabla u\|_{2}^{2}+\left\|\nabla\widetilde{u}_{n}\right\|_{2}^{2}+o_{n}(1),$
$\left\|\nabla v_{n}\right\|_{2}^{2}=$ $\left\|\nabla\left(v+\widetilde{v}_{n}\right)\right\|_{2}^{2}=\|\nabla v\|_{2}^{2}+\left\|\nabla\widetilde{v}_{n}\right\|_{2}^{2}+o_{n}(1),$
$\int_{\mathbb{R}^{N}}\left(|{u}_{n}|^{2}+|{v}_{n}|^{2}\right)\mathrm{d}x=$ $\int_{\mathbb{R}^{N}}\left(|u+\widetilde{u}_{n}|^{2}+|v+\widetilde{v}_{n}|^{2}\right)\mathrm{d}x$ (3.13)
$=$ $\int_{\mathbb{R}^{N}}\left(|{u}|^{2}+|{v}|^{2}\right)\mathrm{d}x+\int_{\mathbb{R}^{N}}\left(|\widetilde{u}_{n}|^{2}+|\widetilde{v}_{n}|^{2}\right)\mathrm{d}x+o_{n}(1)$
and
$\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x=$ $\int_{\mathbb{R}^{N}}|u+\widetilde{u}_{n}|^{r_{1}}|v+\widetilde{v}_{n}|^{r_{2}}\mathrm{d}x$
$=$ $\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x+\int_{\mathbb{R}^{N}}|\widetilde{u}_{n}|^{r_{1}}|\widetilde{v}_{n}|^{r_{2}}\mathrm{d}x+o_{n}(1).$
Hence
$I(u_{n},v_{n})=I(u_{n}\left(x-y_{n}\right),v_{n}\left(x-y_{n}\right))=I(u,v)+I(\widetilde{u}_{n},\widetilde{v}_{n})+o_{n}(1).$ (3.14)
Claim. $\int_{\mathbb{R}^{N}}\left(|\widetilde{u}_{n}|^{2}+|\widetilde{v}_{n}|^{2}\right)\mathrm{d}x\rightarrow 0$ as $n\rightarrow\infty$. That is, $\int_{\mathbb{R}^{N}}\left(|{u}|^{2}+|{v}|^{2}\right)\mathrm{d}x=\rho^{2}$.
In fact, let $\rho^{2}_{1}=\int_{\mathbb{R}^{N}}\left(|{u}|^{2}+|{v}|^{2}\right)\mathrm{d}x>0$. By (3.13), if we get $\rho_{1}=\rho$, the claim follows. Assume $\rho_{1}<\rho$, and define
$\hat{u}_{n}=\left(\frac{\rho^{2}-\rho^{2}_{1}}{\left\|\widetilde{u}_{n}\right\|^{2}_{2}+\left\|\widetilde{v}_{n}\right\|^{2}_{2}}\right)^{1/2}\widetilde{u}_{n},\quad\hat{v}_{n}=\left(\frac{\rho^{2}-\rho^{2}_{1}}{\left\|\widetilde{u}_{n}\right\|^{2}_{2}+\left\|\widetilde{v}_{n}\right\|^{2}_{2}}\right)^{1/2}\widetilde{v}_{n}.$
By (3.14) and Lemma 3.4, it follows that
$I(u_{n},v_{n})=$ $I(u,v)+I(\widetilde{u}_{n},\widetilde{v}_{n})+o_{n}(1)$
$=$ $I(u,v)+I(\hat{u}_{n},\hat{v}_{n})+o_{n}(1)$
$\geq$ $I(u,v)+m\left((\rho^{2}-\rho^{2}_{1})^{1/2}\right)+o_{n}(1).$
Hence, similar to (3.7) and (3.8), we have
$m(\rho)\geq I(u,v)+m\left((\rho^{2}-\rho^{2}_{1})^{{1}/{2}}\right)\geq m\left(\rho_{1}\right)+m\left((\rho^{2}-\rho^{2}_{1})^{1/2}\right)\geq m(\rho).$ (3.15)
Thus $I(u,v)=m\left(\rho_{1}\right)$. That is, $(u,v)$ is global minimizer with respect to $\rho_{1}$. Using Lemma 3.3-(ii), we deduce the strict inequality
$m(\rho)<m\left(\rho_{1}\right)+m\left((\rho^{2}-\rho^{2}_{1})^{1/2}\right),$
which contradicts (3.15). So $\rho_{1}=\rho$, hence we complete the proof of the claim.
Since $\{\left(\widetilde{u}_{n},\widetilde{v}_{n}\right)\}$ is a bounded sequence in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ and using the above claim, it follows from (2.2) and (2) respectively that $\left\|\nabla\widetilde{u}_{n}\right\|_{2}^{2}\rightarrow 0$, $\left\|\nabla\widetilde{v}_{n}\right\|_{2}^{2}\rightarrow 0$ and $\int_{\mathbb{R}^{N}}|\widetilde{u}_{n}|^{r_{1}}|\widetilde{v}_{n}|^{r_{2}}\mathrm{d}x\rightarrow 0$ as $n\rightarrow\infty$. Thus,
$\liminf_{n\rightarrow\infty}I(\widetilde{u}_{n},\widetilde{v}_{n})=\liminf_{n\rightarrow\infty}\frac{1}{2}\left(\|\Delta\widetilde{u}_{n}\|_{2}^{2}+\|\Delta\widetilde{v}_{n}\|_{2}^{2}\right)\geq 0.$ (3.16)
On the other hand, since $\int_{\mathbb{R}^{N}}\left(|{u}|^{2}+|{v}|^{2}\right)\mathrm{d}x=\rho^{2}$, we deduce from (3.14) that
$I(u_{n},v_{n})=I(u,v)+I(\widetilde{u}_{n},\widetilde{v}_{n})+o_{n}(1)\geq m(\rho)+I(\widetilde{u}_{n},\widetilde{v}_{n})+o_{n}(1),$
and so
$\limsup_{n\rightarrow\infty}I(\widetilde{u}_{n},\widetilde{v}_{n})\leq 0.$ (3.17)
From (3.16) and (3.17), we get that
$\left\|\Delta\widetilde{u}_{n}\right\|_{2}^{2}\rightarrow 0,\quad\left\|\Delta\widetilde{v}_{n}\right\|_{2}^{2}\rightarrow 0.$
So, by (3.12), $\left(u_{n}\left(\cdot-y_{n}\right),v_{n}\left(\cdot-y_{n}\right)\right)\rightarrow(u,v)$ in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ as $n\rightarrow\infty$. ∎
With the above alternative argument, we have the following criterion to derive the existence of minimizer for $m(\rho)$.
Proposition 3.6.
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. For $\rho>0$, let $m(\rho)$ and $m^{J}(\rho)$ be given respectively in (1.7) and (1.8). Then $m(\rho)$ can be achieved if $m(\rho)<m^{J}(\rho)$.
Proof.
Suppose first that $m(\rho)<m^{J}(\rho)$. Let $\{(u_{n},v_{n})\}$ be a minimizing sequence of $m(\rho)$. If the sequence $\{(u_{n},v_{n})\}$ vanishes, according to Lions’ lemma [12, Lemma I.1], we have
$\left(u_{n},v_{n}\right)\rightharpoonup$ (0,0) in $(0,0)\ \text{ in }H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}),$
$\left(u_{n},v_{n}\right)\rightarrow$ (0,0) in $(0,0)\ \text{ in }L^{p}\left(\mathbb{R}^{N}\right)\times L^{p}(\mathbb{R}^{N}),\quad\text{ for }2<p<2^{**}.$
Thus
$m(\rho)$ $=I(u_{n},v_{n})+o_{n}(1)$
$=\frac{1}{2}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)-\frac{1}{2}\alpha_{1}\|\nabla u_{n}\|^{2}_{2}-\frac{1}{2}\alpha_{2}\|\nabla v_{n}\|^{2}_{2}+o_{n}(1)$
$=J(u_{n},v_{n})+o_{n}(1)$
$\geq m^{J}(\rho)+o_{n}(1).$
This is a contradiction with $m(\rho)<m^{J}(\rho)$, so $\{(u_{n},v_{n})\}$ does not vanish. Using Theorem 3.5, $\{(u_{n},v_{n})\}$ has a convergent subsequence and $m(\rho)$ is achieved. ∎
3.1 Comparison between $m(\rho)$ and $m^{J}(\rho)$
In this subsection, we will explore what conditions could guarantee $m(\rho)<m^{J}(\rho)$. Now we assume $\alpha_{1}\geq\alpha_{2}$, hence $m^{J}(\rho)=-\frac{\alpha_{1}^{2}}{8}\rho^{2}$. For any $(u,v)\in S_{\rho}$, we have
$H(u,v):=$ $I(u,v)-m^{J}(\rho)$ (3.18)
$=$ $\frac{1}{2}\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}+\frac{1}{2}\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{8}\|v\|_{2}^{2}-\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x$
$=$ $\frac{1}{2}\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\|v\|_{2}^{2}\right)\Big{(}1-2\beta\rho^{r-2}Q(u,v)\Big{)},$
where $Q$ is defined in (1.9). Let us denote
$R:=\sup\,\left\{Q(u,v):(u,v)\in(H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}))\backslash\{(0,0)\}\right\}.$ (3.19)
Next, if $R<\infty$, we denote
$\rho^{*}:=\left(\frac{1}{2\beta R}\right)^{\frac{1}{r-2}}.$ (3.20)
We will see in Proposition 3.8 below that $\rho^{*}$ is the dichotomy parameter for whether $m(\rho)<m^{J}(\rho)$. Note that in the mass-critical case $r=\bar{r}$, Lemma 3.1 tells us that $m(\rho)>-\infty$ only holds for $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. Hence it is important for us whether $\rho^{*}<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$.
Lemma 3.7.
If $r=\bar{r}$, there holds that $\rho^{*}<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$.
Proof.
In view of (3.20), the inequality $\rho^{*}<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ is equivalent to $R>\mathcal{D}_{1}$. To this end, it suffices to show that there exists some $(u_{0},v_{0})$ such that
$Q(u_{0},v_{0})>\mathcal{D}_{1}.$
For that, let $U$ be an extremal of Gagliardo-Nirenberg inequality given in (2.1) with $p=\bar{r}$. Now we take $U_{t}:=U(tx)$, and $u_{t}:=U_{t}$ and $v_{t}:=\sqrt{\frac{r_{2}}{r_{1}}}U_{t}$. Thus
$Q(u_{t},v_{t})=\frac{{(\frac{r_{2}}{r_{1}})}^{\frac{r_{2}}{2}}\int_{\mathbb{R}^{N}}|U_{t}|^{r}\mathrm{d}x}{\left(1+\frac{r_{2}}{r_{1}}\right)^{\frac{r}{2}-1}\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{r_{2}}{r_{1}}\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\frac{r_{2}}{r_{1}}\|U_{t}\|_{2}^{2}\right)\|U_{t}\|_{2}^{r-2}}.$
By (2) and Lemma 2.2, we have
$\left(\frac{r_{2}}{r_{1}}\right)^{\frac{r_{2}}{2}}\|U_{t}\|_{r}^{r}=\mathcal{D}_{1}\left(1+\frac{r_{2}}{r_{1}}\right)^{\frac{r}{2}}\|U_{t}\|_{2}^{r-r\gamma_{r}}\|\Delta U_{t}\|_{2}^{r\gamma_{r}}.$
Consequently, we obtain
$Q(u_{t},v_{t})=\frac{\mathcal{D}_{1}\left(1+\frac{r_{2}}{r_{1}}\right)\|\Delta U_{t}\|_{2}^{r\gamma_{r}}}{\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{r_{2}}{r_{1}}\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\frac{r_{2}}{r_{1}}\|U_{t}\|_{2}^{2}}.$
By scaling, it is easy to get $Q(u_{t},v_{t})>\mathcal{D}_{1}$ as $t$ is large enough. The proof is done. ∎
Proposition 3.8.
  • (i)
    When $R<\infty$, there hold that $m(\rho)=m^{J}(\rho)$ if $\rho\in(0,\rho^{*}]$, and $m(\rho)<m^{J}(\rho)$ if $\rho>\rho^{*}$. Moreover, when $\rho\in(0,\rho^{*})$, $m(\rho)$ is never achieved.
  • (ii)
    When $R=\infty$, we have $m(\rho)<m^{J}(\rho)$ for all $\rho>0$.
Proof.
(i) If $\rho\in(0,\rho^{*}]$, for any $(u,v)\in S_{\rho}$, we obtain
$0\leq 1-2\beta\rho^{r-2}R\leq 1-2\beta\rho^{r-2}Q(u,v).$
Hence it follows from (3.18) that
$\underset{(u,v)\in S_{\rho}}{\inf}H(u,v)\geq 0,$ (3.21)
so $m(\rho)=m^{J}(\rho)$.
Next, we claim that $m(\rho)$ is never achieved for all $\rho\in(0,\rho^{*})$. We have already known that $m(\rho)=-\frac{\alpha_{1}^{2}}{8}\rho^{2}$ for all $\rho\in(0,\rho^{*})$, so $m$ is differentiable and
$m^{\prime}(\rho)=-\frac{\alpha_{1}^{2}}{4}\rho.$ (3.22)
Now we assume by contradiction that $(u,v)$ is a minimizer for $m(\rho)$. Observe that $u\neq 0$ and $v\neq 0$. If otherwise, $(u,v)$ is also a minimizer for $m^{J}(\rho)$, that is a contradiction by Lemma 3.2.
Taking $t>0$ such that $(1+t)\rho<\rho^{*}$, we have
$I((1+t)u,(1+t)v)\geq m((1+t)\rho),$
which together with $I(u,v)=m(\rho)$ implies that
$m^{\prime}(\rho)\rho=$ $\underset{t\to 0^{+}}{\lim}\frac{m((1+t)\rho)-m(\rho)}{t}$ (3.23)
$\leq$ $\underset{t\to 0^{+}}{\lim\inf}\frac{I((1+t)u,(1+t)v)-I(u,v)}{t}$
$=$ $\langle I^{\prime}(u,v),(u,v)\rangle.$
On the other hand, for $10<t<$ 1, we have
$I((1-t)u,(1-t)v)\geq m((1-t)\rho),$
then
$m^{\prime}(\rho)\rho=$ $\underset{t\to 0^{+}}{\lim}\frac{m((1-t)\rho)-m(\rho)}{-t}$ (3.24)
$\geq$ $\underset{t\to 0^{+}}{\lim\sup}\frac{I((1-t)u,(1-t)v)-I(u,v)}{-t}$
$=$ $\langle I^{\prime}(u,v),(u,v)\rangle.$
From (3.22)-(3.24), it follows that
$\langle I^{\prime}(u,v),(u,v)\rangle=-\frac{\alpha_{1}^{2}}{4}\rho^{2}.$ (3.25)
According to the Lagrange multiplier principle, there exists a $\lambda_{\rho}\in\mathbb{R}$ such that
$I^{\prime}(u,v)=\lambda_{\rho}(u,v)\quad\mbox{ in}\;\left(H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})\right)^{-1}.$
This combined with (3.25) gives
$\lambda_{\rho}=-\frac{\alpha_{1}^{2}}{4}.$
Since $(u,v)$ is a solution to (1.1) with $\lambda=-\lambda_{\rho}$, we get
$\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}-\alpha_{1}\|\nabla u\|^{2}_{2}-\alpha_{2}\|\nabla v\|^{2}_{2}+\frac{\alpha_{1}^{2}\rho^{2}}{4}=\beta r\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
Thus
$-\frac{\alpha_{1}^{2}\rho^{2}}{8}=m(\rho)=I(u,v)=-\frac{\alpha_{1}^{2}\rho^{2}}{8}+\left(\frac{r}{2}-1\right)\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
This is a contradiction since $\frac{r}{2}>1$ and $u\neq 0$, $v\neq 0$.
(ii) If $R=\infty$, there must be some $(u,v)\in S_{\rho}$ such that the last term of (3.18) is negative, which means that
$\underset{(u,v)\in S_{\rho}}{\inf}H(u,v)<0,$ (3.26)
that is $m(\rho)<m^{J}(\rho)$. ∎
Proof of Theorem 1.1 completed.
Let $R$ be given as in (3.19), and
$\rho^{*}=\begin{cases}\left(\frac{1}{2\beta R}\right)^{\frac{1}{r-2}}\quad&\mbox{if}\;R<\infty;\\ 0\quad&\mbox{if}\;R=\infty.\end{cases}$
By Lemma 3.7 and Proposition 3.8, we have $m(\rho)<m^{J}(\rho)$ if
$\rho\in\begin{cases}(\rho^{*},\infty)\quad&\mbox{if}\;\;r<\bar{r};\\ \left(\rho^{*},\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r}.\end{cases}$
Using Proposition 3.6, $m(\rho)$ can be achieved. By Proposition 3.8 again, when $\rho<\rho^{*}$, $m(\rho)$ is never achieved. Therefore we can complete the proof. ∎
3.2 Estimates on $R$
According to Theorem 1.1, $m(\rho)$ can be achieved for all $\rho>0$ if $\rho^{*}=0$. Thus, it is very important for us to know what conditions can guarantee $\rho^{*}=0$ or $\rho^{*}>0$. For that, we will discuss the value of $R$ by two cases: $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$.
(i) Case $\alpha_{1}=\alpha_{2}$
Proposition 3.9.
Assume that $\alpha_{1}=\alpha_{2}$. Let $R$ be given in (3.19), then $R$ is finite if and only if (1.10) holds.
Proof.
For simplicity, we set $\alpha:=\alpha_{1}=\alpha_{2}$. By Hölder inequality and Young’s inequality, we can calculate that
$Q(u,v)\leq$ $\frac{\|u\|_{r}^{r_{1}}\|v\|_{r}^{r_{2}}}{\left(\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha}{2}\right)v\right\|_{2}^{2}\right)\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{r}{2}-1}}$ (3.27)
$\leq$ $\frac{\frac{r_{1}}{r}\|u\|_{r}^{r}+\frac{r_{2}}{r}\|v\|_{r}^{r}}{\left(\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha}{2}\right)v\right\|_{2}^{2}\right)\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{r}{2}-1}}$
$\leq$ $\frac{\frac{r_{1}}{r}\|u\|_{r}^{r}+\frac{r_{2}}{r}\|v\|_{r}^{r}}{\left(\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha}{2}\right)v\right\|_{2}^{2}\right)\cdot\frac{1}{2}\left(\|u\|_{2}^{r-2}+\|v\|_{2}^{r-2}\right)}$
$\leq$ $\frac{\frac{r_{1}}{r}\|u\|_{r}^{r}+\frac{r_{2}}{r}\|v\|_{r}^{r}}{\frac{1}{2}\left(\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}+\left\|\left(\Delta+\frac{\alpha}{2}\right)v\right\|_{2}^{2}\|v\|_{2}^{r-2}\right)}$
$\leq$ $\frac{2\max\{r_{1},r_{2}\}}{r}\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}.$
Note that in the last inequality above, we used the basic inequality
$\frac{a+b}{c+d}\leq\max\left\{\frac{a}{c},\frac{b}{d}\right\},\quad\forall a,b,c,d>0.$ (3.28)
Therefore, provided that
$\widetilde{R}:=\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}<\infty,$
then we can get $R<\infty$.
By the scaling transformation $w=u\left(\sqrt{\frac{\alpha}{2}}x\right)$, we have
$\|\Delta w\|_{2}^{2}=\left({\frac{2}{\alpha}}\right)^{\frac{N-4}{2}}\|\Delta u\|_{2}^{2},\quad\|\nabla w\|_{2}^{2}=\left({\frac{2}{\alpha}}\right)^{\frac{N-2}{2}}\|\nabla u\|_{2}^{2},\quad\|w\|_{r}^{r}=\left({\frac{2}{\alpha}}\right)^{\frac{N}{2}}\|u\|_{r}^{r}.$
Consequently,
$\sup_{w\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|w\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha}{2}\right)w\right\|_{2}^{2}\|w\|_{2}^{r-2}}=\left(\frac{2}{\alpha}\right)^{2-\frac{(r-2)N}{4}}\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+1\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}.$ (3.29)
By virtue of [8, Theorem 1.1], we have that (3.29) is finite if and only if (1.10) holds. Combined with (3.27), therefore (1.10) can also guarantee $R<\infty$.
In the following, we shall prove that $R=\infty$ if (1.10) does not hold. In fact, we can take $u=v$ in (1.9), that is
$Q(u,u)=\frac{\int_{\mathbb{R}^{N}}|u|^{r}\mathrm{d}x}{2^{\frac{r}{2}}\left\|\left(\Delta+\frac{\alpha}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}},$
which comes back to the analysis of (3.29). The proof is done. ∎
(ii) Case $\alpha_{1}\neq\alpha_{2}$
Without confusions, we assume $\alpha_{2}<\alpha_{1}$. Similar to the case $\alpha_{1}=\alpha_{2}$, here we still compute the supremum of $Q(u,v)$. By Hölder inequality, Young’s inequality and (3.28), we can get
$Q(u,v)\leq$ $\frac{\|u\|_{r}^{r_{1}}\|v\|_{r}^{r_{2}}}{\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}\right)\left(\|u\|_{2}^{2}+\|v\|_{2}^{2}\right)^{\frac{r}{2}-1}}$
$\leq$ $\frac{\frac{r_{1}}{r}\|u\|_{r}^{r}+\frac{r_{2}}{r}\|v\|_{r}^{r}}{\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}+\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}\right)\cdot\frac{1}{2}\left(\|u\|_{2}^{r-2}+\|v\|_{2}^{r-2}\right)}$
$\leq$ $\frac{\frac{r_{1}}{r}\|u\|_{r}^{r}+\frac{r_{2}}{r}\|v\|_{r}^{r}}{\frac{1}{2}\left(\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}+\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)v\right\|_{2}^{2}\|v\|_{2}^{r-2}\right)}$
$\leq$ $\frac{2\max\{r_{1},r_{2}\}}{r}\max_{j=1,2}\left\{\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha_{j}}{2}\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}\right\}.$
By (3.29), we see that under the condition (1.10), the supremum of $Q$ is finite.
Remark that here we could not deduce what conditions can guarantee $R=\infty$ when $\alpha_{1}\neq\alpha_{2}$, which is an open problem in this paper.
Proof of Theorem 1.3 completed.
The above analysis and Proposition 3.9 can conclude our theorem. ∎
4 Mass-supercritical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r_{1}+r_{2}\in(\bar{r},2^{**})$. We will work in the radial Sobolev space
$H^{2}_{r}(\mathbb{R}^{N}):=\left\{u\in H^{2}(\mathbb{R}^{N}):u\mbox{ is radially symmetric}\right\}$
endowed with the equivalent norm $\|u\|:=\left(\|\Delta u\|^{2}_{2}+\|u\|^{2}_{2}\right)^{\frac{1}{2}}$. For the sake of the compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ with $2<p<2^{**}$, we assume $N\geq 2$. Furthermore, we denote
$S_{\rho}^{r}=\left\{(u,v)\in S_{\rho}:u,\,v\mbox{ are radially symmetric}\right\}.$
From (3.1), we have
$I(u,v)\geq$ $\frac{1}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{1}{2}\mathcal{D}_{2}\rho\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)^{\frac{1}{2}}-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)^{\frac{r\gamma_{r}}{2}}$
=:absent:=:= : $h\left((\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2})^{\frac{1}{2}}\right),$
where the function $h:\mathbb{R}^{+}\rightarrow\mathbb{R}$ is given by
$h(t):=\frac{1}{2}t^{2}-\frac{1}{2}\mathcal{D}_{2}\rho t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}}.$
To understand the geometry structure of the functional $I|_{S_{\rho}}$, we analyze the function $h$.
Lemma 4.1.
Let $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r\in(\bar{r},2^{**})$ and $\beta\rho^{r-2}<c^{*}(N,r)$. Then there exist $0<R_{0}<R_{1}$, such that $h\left(R_{0}\right)=h\left(R_{1}\right)=0$ and $h(t)>0$ if and only if $t\in\left(R_{0},R_{1}\right)$.
Proof.
Setting
$\phi(t):=\frac{1}{2}t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}-1},$
we have that for $t>0$, $h(t)>0$ if and only if $\phi(t)>\frac{1}{2}\mathcal{D}_{2}\rho$. Since $r\gamma_{r}>2$, we see that $\phi(t)$ has a unique critical point $\bar{t}\in(0,+\infty)$, and it is the global maximum point, where
$\bar{t}=\left(\frac{1}{2\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}(r\gamma_{r}-1)}\right)^{\frac{1}{r\gamma_{r}-2}}.$
So, $\max\limits_{t>0}\phi(t)=\phi(\bar{t})=\frac{r\gamma_{r}-2}{2(r\gamma_{r}-1)}\bar{t}$. Clearly, $\phi(+\infty)=h(+\infty)=-\infty$, $\phi(0^{+})=0^{+}$ and $h(0^{+})=0^{-}$, thus
$h(\bar{t})>0\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \phi(\bar{t})>\frac{1}{2}\mathcal{D}_{2}\rho\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \beta\rho^{r-2}<c^{*}(N,r).$
In this case, there exist $0<R_{0}<R_{1}$ such that $h>0$ if and only if $t\in(R_{0},R_{1})$. ∎
Under the assumption $\beta\rho^{r-2}<c^{*}(N,r)$, we denote the set
$A_{\rho}:=\{(u,v)\in S^{r}_{\rho}:\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}<R^{2}_{0}\},$ (4.1)
and it is clear that there exists some $(u_{0},v_{0})\in A_{\rho}$ such that $I(u_{0},v_{0})<0$. Furthermore, we can easily find another point $(u_{1},v_{1})\in S^{r}_{\rho}\backslash A_{\rho}$ such that $I(u_{1},v_{1})<0$. Now, we can define mountain pass set as
$\Gamma_{\rho}:=\{\gamma\in C([0,1],S^{r}_{\rho}):\gamma(0)=(u_{0},v_{0}),\;\gamma(1)=(u_{1},v_{1})\},$
and mountain pass level
$\widetilde{M}(\rho):=\underset{\gamma\in\Gamma_{\rho}}{\inf}\underset{t\in[0,1]}{\max}I(\gamma(t)).$ (4.2)
It can be seen from Lemma 4.1 that $\widetilde{M}(\rho)>0$. In the sequel, we will show that the above level $\widetilde{M}(\rho)$ is a critical value of $I|_{S_{\rho}^{r}}$. We denote the Pohožaev functional
$P(u,v):=2\int_{\mathbb{R}^{N}}(|\Delta u|^{2}+|\Delta v|^{2})\mathrm{d}x-\alpha_{1}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\mathrm{d}x-\alpha_{2}\int_{\mathbb{R}^{N}}|\nabla v|^{2}\mathrm{d}x-2\beta r\gamma_{r}\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
By standard procedures as Jeanjean’s method in [10] and using the augmented functional
$\Psi({(u,v)},s):=$ $I(s\ast(u,v))$
$=$ $\frac{e^{4s}}{2}\int_{\mathbb{R}^{N}}(|\Delta u|^{2}+|\Delta v|^{2})\mathrm{d}x-\frac{\alpha_{1}e^{2s}}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\mathrm{d}x-\frac{\alpha_{2}e^{2s}}{2}\int_{\mathbb{R}^{N}}|\nabla v|^{2}\mathrm{d}x$
$-\beta e^{2r\gamma_{r}s}\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x,\quad\forall(u,v)\in H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N}),\;\;s\in\mathbb{R},$
where $s\ast(u,v):=(s\ast u,s\ast v)$ and $(s\ast u)(x):=e^{\frac{N}{2}s}u\left(e^{s}x\right)$, we can obtain a Palais-Smale sequence approaching Pohožaev manifold for the mountain pass level $\widetilde{M}(\rho)$, here we omit its proof.
Lemma 4.2.
Let $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r\in(\bar{r},2^{**})$ and $\beta\rho^{r-2}<c^{*}(N,r)$. Then there exists a Palais-Smale sequence $\{\left(u_{n},v_{n}\right)\}$ for $I|_{S^{r}_{\rho}}$ at $\widetilde{M}(\rho)$, which satisfies $P\left(u_{n},v_{n}\right)\rightarrow 0$.
The above lemma concludes the existence of a Palais-Smale sequence approaching Pohožaev manifold. We next analyze the compactness of such sequence.
Lemma 4.3.
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r\in(\bar{r},2^{**})$, $N\geq 2$ and $\beta\rho^{r-2}<\min\{c^{*}(N,r),c_{*}(N,r)\}$. Let $(u_{n},v_{n})\subset S^{r}_{\rho}$ be a Palais-Smale sequence for $I|_{S_{\rho}^{r}}$ at some level $c>0$ with $P(u_{n},v_{n})\rightarrow 0$. Then up to a subsequence, $(u_{n},v_{n})\rightarrow(u,v)$ in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$, and $(u,v)$ is a radial solution of problem (1.1)-(1.2) for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
Proof.
From $P(u_{n},v_{n})\rightarrow 0$, (2.4) and $r\in(\bar{r},2^{**})$, we have
$c=$ $I(u_{n},v_{n})+o_{n}(1)$
$=$ $\frac{1}{2}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})-\frac{\alpha_{1}}{2}\|\nabla u_{n}\|^{2}_{2}-\frac{\alpha_{2}}{2}\|\nabla v_{n}\|^{2}_{2}-\beta\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x+o_{n}(1)$
$=$ $\left(\frac{1}{2}-\frac{1}{r\gamma_{r}}\right)(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})-\left(\frac{1}{2}-\frac{1}{2r\gamma_{r}}\right)(\alpha_{1}\|\nabla u_{n}\|^{2}_{2}+\alpha_{2}\|\nabla v_{n}\|^{2}_{2})+o_{n}(1)$
$\geq$ $\left(\frac{1}{2}-\frac{1}{r\gamma_{r}}\right)(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})-\left(\frac{1}{2}-\frac{1}{2r\gamma_{r}}\right)\mathcal{D}_{2}\rho(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})^{\frac{1}{2}}+o_{n}(1).$
Combined with $(u_{n},v_{n})\subset S^{r}_{\rho}$, we can deduce that $\{(u_{n},v_{n})\}$ is bounded in $H^{2}_{r}(\mathbb{R}^{N})\times H^{2}_{r}(\mathbb{R}^{N})$.
By the compactness of the embedding $H^{2}_{r}(\mathbb{R}^{N})\hookrightarrow L^{p}\left(\mathbb{R}^{N}\right)$ for $2<p<2^{**}$, there exists a $(u,v)\in H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$ such that up to a subsequence,
$\left(u_{n},v_{n}\right)\rightharpoonup$ $(u,v)\quad\text{ in }H^{2}_{r}(\mathbb{R}^{N})\times H^{2}_{r}(\mathbb{R}^{N}),$ (4.3)
$\left(u_{n},v_{n}\right)\rightarrow$ $(u,v)\quad\text{ in }L^{p}\left(\mathbb{R}^{N}\right)\times L^{p}\left(\mathbb{R}^{N}\right),\text{ for }2<p<2^{**},$
$\left(u_{n},v_{n}\right)\rightarrow$ $(u,v)\quad\text{ a.e.\leavevmode\nobreak\ in }\mathbb{R}^{N}\times\mathbb{R}^{N}.$
Since $I^{\prime}|_{S^{r}_{\rho}}(u_{n},v_{n})\rightarrow 0$, by the Lagrange multipliers rule, we get that there exists a sequence $\{\lambda_{n}\}\subset\mathbb{R}$ such that
$I^{\prime}(u_{n},v_{n})+\lambda_{n}(u_{n},v_{n})\to 0\quad\mbox{in }\left(H^{2}_{r}(\mathbb{R}^{N})\times H^{2}_{r}(\mathbb{R}^{N})\right)^{*}.$ (4.4)
Multiplying (4.4) by $(u_{n},v_{n})$, we have
$\lambda_{n}\rho^{2}=$ $-(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})+\alpha_{1}\|\nabla u_{n}\|^{2}_{2}+\alpha_{2}\|\nabla v_{n}\|^{2}_{2}+\beta r\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x+o_{n}(1).$ (4.5)
Thus by (2), we deduce that $\{\lambda_{n}\}$ is bounded, and hence up to a subsequence $\lambda_{n}\rightarrow\lambda$ for some $\lambda\in\mathbb{R}$.
Next, we claim that $\lambda>0$, $u\neq 0$ and $v\neq 0$. In fact, by (4.5), $\lambda_{n}\rightarrow\lambda$, $P(u_{n},v_{n})\rightarrow 0$ and $r\in(\bar{r},2^{**})$, we have
$\lambda\rho^{2}=\lim_{n\rightarrow\infty}\left(\frac{\alpha_{1}}{2}\|\nabla u_{n}\|^{2}_{2}+\frac{\alpha_{2}}{2}\|\nabla v_{n}\|^{2}_{2}\right)+\frac{\beta}{2}\left(N-r\frac{N-4}{2}\right)\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x\geq 0.$ (4.6)
If $\lambda=0$, it follows from (4.6) that
$\lim_{n\rightarrow\infty}\|\Delta u_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\Delta v_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\nabla u_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\nabla v_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x=0.$
So $I(u_{n},v_{n})\rightarrow 0$, which contradicts our assumption that $I(u_{n},v_{n})\rightarrow c>0$. Thus, there holds $\lambda>0$.
If $u=0$ or $v=0$, we deduce from (4.3) that
$\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x=0.$ (4.7)
With (4.7) and using $P(u_{n},v_{n})\rightarrow 0$ again,
$0<c=$ $\lim_{n\rightarrow\infty}I(u_{n},v_{n})=\lim_{n\rightarrow\infty}\left(\frac{1}{2}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})-\frac{\alpha_{1}}{2}\|\nabla u_{n}\|^{2}_{2}-\frac{\alpha_{2}}{2}\|\nabla v_{n}\|^{2}_{2}\right)$
$=$ $-\frac{1}{2}\lim_{n\rightarrow\infty}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)\leq 0,$
which is impossible. Thus, $u\neq 0$ and $v\neq 0$.
Claim. $\liminf\limits_{n\rightarrow\infty}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)$ has a positive lower bound and $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
According to $P(u_{n},v_{n})\rightarrow 0$, we deduce that
$0<c=I(u_{n},v_{n})+o_{n}(1)=-\frac{1}{2}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})+\beta(r\gamma_{r}-1)\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x+o_{n}(1).$
This combined with (2) shows that
$\frac{1}{2}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})\leq$ $\beta(r\gamma_{r}-1)\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x+o_{n}(1)$ (4.8)
$\leq$ $\beta(r\gamma_{r}-1)\mathcal{D}_{1}\rho^{r(1-\gamma_{r})}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)^{\frac{r\gamma_{r}}{2}}+o_{n}(1).$
From $r\in(\bar{r},2^{**})$, we obtain
$\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})\geq\left(\frac{1}{2\beta(r\gamma_{r}-1)\mathcal{D}_{1}\rho^{r(1-\gamma_{r})}}\right)^{\frac{2}{r\gamma_{r}-2}}.$ (4.9)
Next, we prove that $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. Firstly we get from $P\left(u_{n},v_{n}\right)\rightarrow 0$ that
$\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x\leq\frac{1}{\beta r\gamma_{r}}\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}).$ (4.10)
In the sequel, we prove the claim by two cases.
Case. $\frac{1}{2}<\gamma_{r}<1$. Using $P\left(u_{n},v_{n}\right)\rightarrow 0$, (4.5), (4.9) and (4.10), we have
$\lambda\rho^{2}=$ $\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})+\beta r(1-2\gamma_{r})\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x$
$\geq$ $\frac{1-\gamma_{r}}{\gamma_{r}}\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})$
$\geq$ $\frac{1-\gamma_{r}}{\gamma_{r}}\left(\frac{1}{2\beta(r\gamma_{r}-1)\mathcal{D}_{1}\rho^{r(1-\gamma_{r})}}\right)^{\frac{2}{r\gamma_{r}-2}}.$
Since $\beta\rho^{r-2}<c_{*}(N,r)$, thus $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
Case. $0<\gamma_{r}\leq\frac{1}{2}$. By $P\left(u_{n},v_{n}\right)\rightarrow 0$, (4.5), (4.8) and (4.9), we have
$\lambda\rho^{2}=$ $\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})+\beta r(1-2\gamma_{r})\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x$
$\geq$ $\frac{r-2}{2(r\gamma_{r}-1)}\liminf_{n\rightarrow\infty}(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2})$
$\geq$ $\frac{r-2}{2(r\gamma_{r}-1)}\left(\frac{1}{2\beta(r\gamma_{r}-1)\mathcal{D}_{1}\rho^{r(1-\gamma_{r})}}\right)^{\frac{2}{r\gamma_{r}-2}}.$
Since $\beta\rho^{r-2}<c_{*}(N,r)$, thus $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. We complete the proof of the claim.
Since $u_{n}\rightharpoonup u\neq 0$ and $v_{n}\rightharpoonup v\neq 0$ weakly in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$ and by (4.4), we have
$I^{\prime}(u,v)+\lambda(u,v)=0\quad\mbox{in }\;\left(H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})\right)^{*}.$ (4.11)
Let $(u_{n}-u,v_{n}-v)$ multiply (4.4) and (4.11). By $\lambda_{n}\rightarrow\lambda$, we obtain
$(I^{\prime}(u_{n},v_{n})-I^{\prime}(u,v))[u_{n}-u,v_{n}-v]+\lambda\int_{\mathbb{R}^{N}}(|u_{n}-u|^{2}+|v_{n}-v|^{2})\mathrm{d}x=o_{n}(1).$
Since $(u_{n},v_{n})\rightarrow(u,v)$ in $L^{p}(\mathbb{R}^{N})\times L^{p}(\mathbb{R}^{N})$ for $2<p<2^{**}$ and using (2.4), we deduce that
$\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(u_{n}-u)\|^{2}_{2}+\lambda(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})$ (4.12)
$=$ $\alpha_{1}\|\nabla(u_{n}-u)\|^{2}_{2}+\alpha_{2}\|\nabla(v_{n}-v)\|^{2}_{2}+o_{n}(1)$
$\leq$ $\max\{\alpha_{1},\alpha_{2}\}\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}+o_{n}(1).$
Assume that $\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\geq\delta$ and $\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2}\geq\delta$ for some $\delta>0$. By (4.12), we have
$2\sqrt{\lambda}\leq$ $\frac{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}+\lambda\frac{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}$
$\leq$ $\max\{\alpha_{1},\alpha_{2}\}+o_{n}(1),$
which contradicts $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. So $\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\rightarrow 0$ or $\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2}\rightarrow 0$. In both cases, using again (4.12) and $\lambda>0$, we can prove that
$\|\Delta(u_{n}-u)\|^{2}_{2},\leavevmode\nobreak\ \|\Delta(v_{n}-v)\|^{2}_{2},\leavevmode\nobreak\ \|u_{n}-u\|^{2}_{2},\leavevmode\nobreak\ \|v_{n}-v\|^{2}_{2}\rightarrow 0\quad\text{ as }n\rightarrow\infty.$
Thus, $(u_{n},v_{n})\rightarrow(u,v)$ strongly in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$. ∎
Proof of Theorem 1.5 completed.
Under $\beta\rho^{r-2}<c^{*}(N,r)$, let $\widetilde{M}(\rho)$ be given in (4.2). By Lemma 4.2, we obtain a Palais-Smale sequence $(u_{n},v_{n})$ for $I|_{S^{r}_{\rho}}$ at the level $\widetilde{M}(\rho)$, which satisfies $P\left(u_{n},v_{n}\right)\rightarrow 0$. Using Lemma 4.2, if in addition $\beta\rho^{r-2}<c_{*}(N,r)$, we can obtain a radial solution $(u,v)$ for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. ∎